OpenStudy (anonymous):

What is the value of the expression? 19.2 + (–7 + 3.8) A.16 B.24 C.29 D.30

jigglypuff314 (jigglypuff314):

order of operations: PEMDAS so do the stuff in the Parenthesis first. so 19.2 + (–7 + 3.8) = 19.2 + (-3.2) = 19.2 - 3.2 = ...

OpenStudy (anonymous):

A.

OpenStudy (the_fizicx99):

Order of Operations. 19.2 + (–7 + 3.8) 19.2 + (-3.2) = 16 So, A.